The area of a rectangle can be expressed as x^2+7x-30
What is the legnth and width of the rectangle?
Can you please show steps :)

Answer by sachi(548)

x^2+7x-30=x^2+(10-3)x-30=x^2+10x-3x-30=x(x+10)-3(x+10)=(x+10)(x-3)
so length=(x+10)
& width=(x-3)
ans